May you pls help me with this question
"In 9 years time,the mother will be twice as old as her son,she was 4 times older than her son after 3 years.find their age?
Thank you tutor.
kaaliendjala@gmail.com

Found 2 solutions by Theo, ikleyn:
Answer by Theo(13342)

if i interpret the question correctly, it looks like we have a very young mother.
the answer, as far as i can tell, is the age of the mother today is 9 years old and the son was just born so the age of the son is 0 years old.
in 9 years, the mother will be 18 and the son will be 9.
this means the mother is twice as old as the son.
in 3 years, the mother will be 12 and the son will be 3.
this means the mother is 4 times as old as the son.
unless i misunderstood the problem, this is the answer i get.
the problem as far as i can see is this.
statement 1:
In 9 years time,the mother will be twice as old as her son.
if we allow m to represent her age now, and s to represent her son's age now, then the equation we get from this statement is this:

m+9 = 2*(s+9)

statement 2:
she was 4 times older than her son after 3 years.

i interpreted this to mean that she was 4 times as old as her son in 3 years.
the equation for this would be:

m+4 = 4*(s+3)

if the interpretations are correct, then the mathematics works out as follows:

you have 2 equations that need to be solved simultaneously.
they are:

m+9 = 2*(s+9)
m+3 = 4*(s+3)

simplify these equations to get:

m+9 = 2*s+18
m+3 = 4*s+12

there are several ways to solve this, but the answer becomes s = 0 and m = 9
to confirm that this is the correct answer, solve both equations by substituting 0 for s and 9 for m.

the first equation of:
m+9 = 2*s+18 becomes:
9+9 = 2*0+18 which becomes:
18 = 18 which confirms that m = 9 and s = 0 works for the first equation.

the second equation of:
m+3 = 4*s+12 becomes:
9+3 4*0+12 which becomes:
12 = 12 which confirms that m = 9 and s = 0 works for the second equation.

since m = 9 and s = 0 works for both equations, that is the common solution to both equations.

it might look strange (a 9 year old mother would be very rare), but the mathematics comes out, assuming the interpretation of the problem is correct.





Answer by ikleyn(53751)

.
May you pls help me with this question
"In 9 years time, the mother will be twice as old as her son, she was 4 times older than her son after 3 years. find their age?
Thank you tutor.
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~


This post violates the rules of  English grammar in part of sequence of tenses :  the words and the times are not consistent.
So, if to read/understand it literally as it is written,  it makes no sense and is not an object for discussions.

So,  in this form,  this written  " problem "  is nonsense.

There are two ways to react on it :   either throw to a garbage bin - - - or edit it from scratch to make it correctly.


Also,  for your information,  this form

    "she was 4 times older than her son after 3 years."

is not used in  Math problems written in  English,  since in  English,  unlike many other languages,
this form creates ambiguity unacceptable in  Math problems.


It is why the other tutor obtained absurdist answer in his post.